Neural Computation (ISSN 0899-7667) is published quarterly by The MIT Press, Cambridge, Massachusetts.

Please see “Information for Contributors” at the back of this issue for information regarding submissions.

Formatted with LaTEX by Archetype Publishing Inc., P.O. Box 6567, Champaign, IL 61821. Printed in the U.S.A.

Subscription Information

Business Offices: Send subscriptions and address changes to MIT Press Journals, 55 Hayward Street, Cambridge, MA 02142. Phone: (617) 253-2889.

Subscription Rates: $45.00 for individuals, $90.00 for institutions, $35.00 for students. Subscribers outside the United States and Canada add $9.00 for surface mail or $17.00 for airmail postage. Single copies of current issues: $25.00. Checks must be drawn on U.S. banks in U.S. funds. To be honored free, claims for missing issues must be made immediately upon receipt of the next published issue.

Permission Information

The copyright owner has authorized the Copyright Clearance Center (CCC), Transactional Reporting Service to grant permission to photocopy articles from this journal for personal use or use internal to an organization, provided that the user reports the fee-code for usage (0899-7667/89 $5.00) and pays the fee of $5.00 per article-copy, to the CCC. Pay fees directly, at the time of reporting, to CCC, Transactional Reporting Service, 27 Congress Street, Salem, MA 01970. For organizations that have been granted a photocopy license with CCC, a separate system of payment has been arranged. Address all other inquiries to the Journals Permissions Manager, MIT Press Journals, 55 Hayward Street, Cambridge, MA 02142.

Postmaster: Send address changes to Neural Computation, 55 Hayward Street, Cambridge, MA 02142. Second class postage pending at Boston, MA, and at additional post offices.

Advertising: Inquiries may be addressed to the Advertising Manager, MIT Press Journals, 55 Hayward Street, Cambridge, MA, 02142. (617) 253-2866.

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989 by the Massachusetts Institute of Technology

Cover Illustration: Neural networks for curve detection. See Zucker, Dobbins, and Iverson, page 75.